OpenStudy (anonymous):

S = 4Lw + 4wh + 3hL solve for h. Please help explain in steps.

OpenStudy (greatlife44):

\[\frac{ S-4Lw }{ 4w+3L} = h\]

OpenStudy (greatlife44):

I had written out a detailed explanation but somehow my equation bar froze and my answer was deleted.

OpenStudy (greatlife44):



OpenStudy (greatlife44):

we manipulate the distributive law which says the following: \[a(b+c) = a*b + a*c \]

OpenStudy (greatlife44):

\[h(4w+3l) = h*4w + h*3L \] do you see how this is just one way of writing the same thing? because h is on one side we can isolate it.

OpenStudy (greatlife44):

@mydearmegan

OpenStudy (anonymous):

@greatlife44 thank you so much! I understand it! I think I was overthinking the whole thing because now it seems so simple. You rock.

OpenStudy (greatlife44):

@mydearmegan look at the number of terms you have then see what variables you need to isolate. you isolate by doing the opposite operation